express the logarithm in terms of log2 M and log2 N: log2 cube root of M^2N
Answer by fractalier(1829)

log_2 (cbrt (M^2N)) =
(1/3)*(log_2 (M^2N)) =
(1/3)*(log_2 M^2 + log_2 N) =
(1/3)*(2*log_2 M + log_2 N) =
(2/3)*log_2 M + (1/3)*log_2 N